OpenStudy (anonymous):

Joey is going to paint his room. His ceiling is 8ft high. The room is 12ft wide and 15ft long. What is the total area of the walls? (Calculate as if there were no windows or doors.) A. 1440 ft3 B. 1440 ft2 C. 432 ft3 D. 432 ft2

OpenStudy (anonymous):

I think B, I just need someone to check it please

OpenStudy (anonymous):

Can someone please tell me if this is right, ill fan and medal

OpenStudy (ybarrap):



OpenStudy (anonymous):

I know that, I just needed to know if my answer was right

OpenStudy (ybarrap):

This is the situation. Area is the sum of all sides: There are 2 walls that are 8 ft high and 12 ft wide: 2 * 8*12 Two walls that are 8 ft high and 15 ft long: 2 * 8 * 15 One ceiling that is 15 ft long and 12 ft wide: 15 * 12 So to get total area we sum all these to get $$ \text{Area = }2\times8\times 12+2\times 8\times 15+15\times 12 $$ Is that what you did?